# Entropy production of the contact model

### v1

@@ Abstract
We propose an expression for the production of entropy for system described by a stochastic dynamics which is appropriate for the case where the reverse transition rate vanishes but the forward transition is nonzero. The expression is positive defined and based on the inequality $x\ln x-(x-1)\geq 0$. The corresponding entropy flux is linear in the probability distribution allowing its calculation as an average. The expression is applied to the one-dimensional contact process at the stationary state. We found that the rate of entropy production per site is finite with a singularity at the critical point with diverging slope.
@@ I Introduction
Let us consider a stochastic dynamics in discrete time and let $P_{i}$ be the probability distribution of states $i$, and $W_{ij}$ be the transition probability rate from state $j$ to state $i$. The equation governing the time evolution of $P_{i}$ is assumed to be given by a master equation kampen1981 ; tome2015L
@@
which is always nonnegative, reflecting the main feature of the production of entropy. This is the usual expression for the rate of entropy production and has been extensively studied and applied luo1984 ; mou1986 ; lebowitz1999 ; maes2003 ; crochik2005 ; seifert2005 ; tome2006 ; zia2007 ; andrieux2007 ; harris2007 ; gaveau2009 ; esposito2009 ; szabo2010 ; tome2010 ; hinrichsen2011 ; benavraham2011 ; tome2012 ; esposito2012 ; zeraati2012 ; tome2015 ; tome2018 ; busiello2020 ; tome2023 . In particular it has been used to determined the rate of entropy production of irreversible stochastic lattice models crochik2005 ; tome2012 ; tome2023 at the nonequilibrium stationary state.
The expression (2) however breaks down when one transition rate vanishes while the other remains nonzero, that is, when a transition has no reverse. There are many examples of stochastic systems where this happens, the most striking examples being the systems with absorbing states such as the contact process tome2015L ; marro1999 ; hinrichsen2000 ; henkel2008 . In this model, particles residing on a lattice are annihilated and created according to the following rules. A particle is annihilated spontaneously, and is created in a empty site if this site has at least one neighboring site occupied. Thus the transition corresponding to the annihilation of an isolated particle has no reverse, and the application of the formula (2) would result in a divergent rate of entropy production.
There has been some attempts to overcome this problem. These attempts focused mainly on the introduction of an effective backward transitions while keeping formula (2) unchanged benavraham2011 ; zeraati2012 . To determine the rate of entropy production when the reverse transition is absent, that is, when $W_{ij}>0$ but $W_{ji}=0$, we propose the following expression
@@ IV Contact model
The results above indicates that both the production of entropy $\psi$ and its derivative $\Gamma=d\psi/dp$ have a singular behavior at the critical point that is described by a power law with exponents $b$ and $a$. We point out that the numerical value of the exponent $b$ for the production of entropy is numerically equal to the exponent $\beta$ associated to the order parameter $\rho$ of the one-dimensional contact model which is $\beta=0.2765$ marro1999 .
@@
V Conclusion
We have proposed an expression for the rate of entropy production for the case in which the reverse transition rate vanishes but the forward transition rate is nonzero. This expression is positive defined and the corresponding entropy flux is linear in the probability distribution allowing its calculation as an average over this distribution. This expression allows us to determine the rate of entropy production for stochastic dynamic systems whose some of the transitions do not have the reverse. This is the case of models with with absorbing states such as the contact process.

### v2

@@ Abstract
We propose an expression for the production of entropy for system described by a stochastic dynamics which is appropriate for the case where the reverse transition rate vanishes but the forward transition is nonzero. The expression is positive definite and based on the inequality $x\ln x-(x-1)\geq 0$. The corresponding entropy flux is linear in the probability distribution allowing its calculation as an average. The expression is applied to the one-dimensional contact process at the stationary state. We found that the rate of entropy production per site is finite with a singularity at the critical point with diverging slope.
@@ I Introduction
Let us consider a stochastic dynamics in continuous time and discrete space of states and let $P_{i}$ be the probability distribution of states $i$, and $W_{ij}$ be the transition probability rate from state $j$ to state $i$. The equation governing the time evolution of $P_{i}$ is assumed to be given by a master equation kampen1981 ; tome2015L
@@
which is always nonnegative, reflecting the main feature of the production of entropy. This is the usual expression for the rate of entropy production and has been extensively studied and applied luo1984 ; mou1986 ; lebowitz1999 ; maes2003 ; seifert2005 ; tome2006 ; zia2007 ; andrieux2007 ; harris2007 ; gaveau2009 ; esposito2009 ; szabo2010 ; tome2010 ; hinrichsen2011 ; esposito2012 ; tome2015 ; tome2018 . In particular it has been used to determined the rate of entropy production of irreversible stochastic lattice models crochik2005 ; tome2012 ; noa2019 ; tome2023 ; hawthorne2023 at the nonequilibrium stationary state.
The expression (2) however breaks down when one transition rate vanishes while the other remains nonzero, that is, when a transition has no reverse. There are many examples of stochastic systems where this happens, the most striking examples being the systems with absorbing states such as the contact process marro1999 ; hinrichsen2000 ; henkel2008 . In this model, particles residing on a lattice are annihilated and created according to the following rules. A particle is annihilated spontaneously, and is created in a empty site if this site has at least one neighboring site occupied. Thus the transition corresponding to the annihilation of an isolated particle has no reverse, and the application of the formula (2) would result in a divergent rate of entropy production.
It should be pointed out that the Schnakenberg formula has a straight relationship with the Jarzynski equality. Thus the breakdown of this formula results in the inapplicability of the Jarzynski inequality when the probability of the backward probability vanishes wereas the forward probability remains finite murashita2014 .
This problem has been addressed by several authors and there have been some attempts to overcome it murashita2014 ; ohkubo2009 ; benavraham2011 ; zeraati2012 ; rahav2014 ; saha2016 ; pal2017 ; busiello2020 ; pal2021 ; pal2021a . Some of these attempts focused mainly on the introduction of an effective backward transitions while keeping formula (2) unchanged. A distinguishing formula was proposed by Busiello, Gupta, and Maritan busiello2020 and will be commented further on. Here we take a different route. To determine the rate of entropy production when the reverse transition is absent, that is, when $W_{ij}>0$ but $W_{ji}=0$, we propose the following expression
@@ IV Contact model
The results above indicates that both the production of entropy $\psi$ and its derivative $\Gamma=d\psi/dp$ have a singular behavior at the critical point that is described by a power law with exponents $b$ and $a$. We point out that the numerical value of the exponent $b$ for the production of entropy is numerically equal to the exponent $\beta$ associated to the order parameter $\rho$ of the one-dimensional contact model which is $\beta=0.2765$ marro1999 . This result comes from the fact that the entropy production is dominated by $P(010)$ which has a critical behavior similar to the density of particles which is the order parameter.
@@
V Unidirectional transition rates
Let us consider that the states 1 and 2 are connected by a unidirectional transition rate $W_{21}>0$ from 1 to 2. The contribution of this transition rate to $dP_{1}/dt$ is $-W_{21}P_{1}$ and to $dP_{2}/dt$ is $W_{21}P_{1}$, and the contribubtion to $dS/dt$ is
$W_{21}P_{1}\ln\frac{P_{1}}{P_{2}}.$ (33)
Let us consider the case of a system such that at the stationary state the difference $P_{1}-P_{2}$ is small. In this case the expression above becomes
$-W_{21}(P_{2}-P_{1}).$
At the stationary state $P_{1}$ and $P_{2}$ does not depend on time and $P_{2}\geq P_{1}$, and this contribution to $dS/dt$ is negative. Since at the stationary state $dS/dt$ vanishes then there should be positive contributions coming from the bidirectional transitions. If the transitions are all unidirectional then at the stationary state $P_{1}=P_{2}$ and all probabilities are equal. Taking into account that $dS/dt$ is equal to the entropy production $\Pi$ minus the flux of entropy $\Psi$ then the term above should be a part of the flux of entropy, that is, the contribution to $\Psi$ coming from the unidirectional transition is
$\Psi_{21}=W_{21}(P_{2}-P_{1}).$ (34)
Summing this term with (33), we find the contribution of the unidirectional transition to the rate of entropy production
$\Pi_{21}=W_{21}P_{1}\ln\frac{P_{1}}{P_{2}}+W_{21}(P_{2}-P_{1}).$ (35)
which turns out to be positive definite, and is our proposed expression (4).
The above reasoning does not show that the form (34) is necessary but shows that it is sufficient. In fact any other form $F(P_{2},P_{1})$ such that $F\geq 0$ for $P_{2}\geq P_{1}$ will serve as long as
$P_{1}\ln\frac{P_{1}}{P_{2}}+F(P_{2},P_{1})\geq 0,$ (36)
for any $P_{1}$ and $P_{2}$. Writing $F(P_{2},P_{1})=P_{1}f(P_{2}/P_{1})$ this condition becomes
$f(x)\geq\ln x,$ (37)
supplemented by $f(1)=0$. Our choice is $f(x)=x-1$. The choice made by Busiello, Gupta, and Maritan busiello2020 is
$\Psi_{21}=W_{21}P_{1}\ln\frac{P_{2}}{P_{1}}.$ (38)
and corresponds to $f(x)=\ln x$. But this expression for $\Psi_{21}$ makes the corresponding entropy production $\Pi_{21}=0$ to vanish identically.
An essential property required for the production of entropy is that it vanishes at thermodynamic equilibrium. As the production of entropy is composed of a sum of terms that are each one nonnegative then each term should vanish. In the case of the Schnakenberg formula (2), the vanishing of each term implies that the transition rates should fulfill the condition
$W_{ij}P_{j}^{0}=W_{ji}P_{i}^{0},$ (39)
where $P_{i}^{0}$ is the equilibrium distribution. In the case of formula (4), the condition is
$P_{j}^{0}=P_{i}^{0}.$ (40)
Therefore if both types of transitions are present, unidirectional and bidirectional, then the thermodynamic equilibrium is possible if $W_{ij}=W_{ji}$.
We show next that the production of entropy is a minimum at the equilibrium. To this end we write $P_{i}=q_{i}+\delta p_{i}$ and replace in
$\Pi_{ij}=W_{ij}[q_{j}\ln\frac{q_{j}}{q_{i}}-(q_{j}-q_{i})].$ (41)
Up to second order in $\delta p_{i}$, we find
$\delta\Pi_{ij}=W_{ij}[\delta p_{j}\ln\frac{q_{j}}{q_{i}}-\delta p_{i}(\frac{q_{j}}{q_{i}}-1)]+\frac{1}{2}W_{ij}q_{j}(\frac{\delta p_{j}}{q_{j}}-\frac{\delta p_{i}}{q_{i}})^{2}.$ (42)
If $q_{i}=P_{i}^{0}$ the terms linear in $\delta p_{i}$ vanish and we are left with
$\delta\Pi_{ij}=\frac{W_{ij}}{2P_{i}^{0}}(\delta p_{j}-\delta p_{i})^{2},$ (43)
which shows that the entropy production is a minimum at equilibirum. The value of $\Pi_{ij}$ at equilibrium also vanish as expected.
We remark that the additivity of the entropy flux follows immediately from the expression (22) and is implicit is our numerical calculation. Notice that the plots in figure 1 shows the entropy flux per site of the lattice.
VI Conclusion
We have proposed an expression for the rate of entropy production for the case in which the reverse transition rate vanishes but the forward transition rate is nonzero. This expression is positive definite and the corresponding entropy flux is linear in the probability distribution allowing its calculation as an average over this distribution. This expression allows us to determine the rate of entropy production for stochastic dynamic systems whose some of the transitions do not have the reverse. This is the case of models with with absorbing states such as the contact process.